OpenStudy (anonymous):

Given f(x)=x^2-4x-3 and g(x)=x+3 over 4. Solve for f(g(x)) when x=9 A.-6 B. 3 C.6 D.9

OpenStudy (valpey):

\[g(x)=\frac{x+3}{4}\]\[g(9)=\frac{9+3}{4}= \frac{12}{4}=3\]

OpenStudy (anonymous):

A

OpenStudy (valpey):

\[f(g(9))= f(3)\]\[f(x)=x^2-4x-3\]\[f(3)=3^2-4*3-3= 9-12-3\]